the dish together with 11 balls weighs 1.6kg, the weight of one ball is 100g. how much does an empty container weigh?


1
Expert's answer
2021-09-14T14:15:56-0400

"M+11m=1.6\to M=1.6-11m=1.6-11\cdot0.1=0.5\ (kg)"